Fermi National Laboratory
 

Fermilab Shutdown Report for February 24, 2006 - March 3, 2006

 

LINAC

March 1 - 3
Install New 750 KeV Beam Stop
Interlock New Beam Stop
Started MuCool Magnet Move
Continue I- Anode Supply Upgrade
Replace water fittings in Klystron Modulator De-Qing circuit
Clean LRF1, LRF2 & LRF3 Modulator and Rectifiers
Replaced Collimator between LE LINAC Tanks 1 & 2
Install new filters and other maintenance on Water Systems
LE LINAC LLRF System Characterization Started on LRF4

March 6 & 7
Finish MuCool Magnet Move
Continue I- Source Anode Supply Upgrade
Clean LRF4 & LRF5 Modulator & Rectifier
LE LINAC LLRF System Characterization Continued on LRF4 & LRF5

March 8 (L.E. LINAC Turn On)
LE LINAC (Stations 1-3) Turn ON

March 9 & 10
NTF Patient Treatment
Start maintenance on H.E. Klystrons Systems
Investigate Q3BULK Supply Failure
LE LINAC LLRF System Characterization on LRF5

Booster
Surveys, ALARAs, Lockouts, HAs completed
No hitches getting workers in and started!!

MI-8
Completed MI-8 tunnel trench work – sump
Shield wall removal about 50% complete
Just starting MI-8 excavation

400 MeV
400 MeV Power Supply work going well
ORBMP and Girder testing progressing

Pbar
Completed rotating BPMs

Main Injector
Staging equipment
Began installation of MI8 collimator
Began installation of MI10 kicker cooling system

 

 

 

Recycler

 

 

Tevatron
Jobs around the TeV

Shim smart bolts on last 244 dipoles
Roll measurements dipoles and quadrupoles 2 sectors done will finish this weekend.
Install water levels all quads in TeV B,E done, C needs electrical checking A,D,F underway
Level run around TeV check elevations start week 4
As found low beta quads and separators B0 and D0. D0 done B0 underway
As found 2 sectors with laser tracker start week 7
Upgrade HLS low beta quads B0 and D0
Cryo and Electrical maintenance

TeV Vacuum work
Install separators A49, B48, A17
There will be two shifts of leak checking
Replace F47-2 dipole
Fix leaks A-3 house and replace A38-1 and un roll A39-1
Fix cold leak E-2
Investigate leaks B-4
D-1 and E-3 on hold
Install TEL 2
Crystal collimator, IPM work at E-0

 

 

SwitchYard
No Report

NuMI
The target chase air-cooling skid has been disconnected
The R-blocks have been removed
Investigating target hall air leaks

MiniBooNE
No Report

Collider Detector at Fermilab
They had finished their cosmic run
The COT is off gas
Their detector has been locked off

D-Zero
Their shutdown is driven by their level 0 install

Alignment
Main Injector
Completed “As Founds” of the wide aperture quadrupoles (quads)
MI-8
Staked 4 stands and BPM
Tevatron
Completed "As Founds" of the low beta quads and separators at D0
Marked locations of Deep Rod monuments at sectors A, E, and F straight sections
SwitchYard
Measured 4 quads at sector F49 in the Transfer Hall
Completed “As Founds” of quads and dipoles for sectors F46, F47, of the P3 line
MiniBooNE
Found fiducials for the MiniBooNE Horn
D-Zero
Completed "As Founds" of the Roman Pots at sector A48

Mechanical
Janitorial work in the tunnels has begun

Cryogenics
Cryo Operations
Tevatron cryogenic system shutdown went smoothly with helium inventory transferred to CHL's LHe storage Dewar and GHe Tank Farm.  CHL NRL plant is still online though at reduced capacity for lower LN2 demand.
Tevatron at LN2 temperature and stable with magnet cryostat vacuum problems addressed well by MSD/Vacuum Group.
F4 and A3 are at room temperature and E2 is slowly warming up to room temperature.
Failure of F47-2 magnet cryostat vacuum was found to be caused by a Kautzky relief valve that failed in closed position, thus resulting in a trapped volume condition leading to rupture.   All 1-phase Kautzky relief valves were checked for proper operation at F4, A3, and A1 and all were found functioning properly.  Failure diagnosis is in progress with problem resolution plan to be developed.
ACNET application programs relating to FRIG operations are being tested on Linux consoles.
CHL shift work (24/7) operations will cease on next Friday after purification of GHe Tank Farm and shutdown of NRL for maintenance.

CHL Maintenance
General facility maintenance was started.
CHL Helium Compressor “A” cooling tower installation is continuing with anticipated completion by month end.

FRIG Maintenance
Ring wide overhauls of wet expanders were started.
Selected wet expander flywheel replacements are in progress.
Leak checking for vacuum leaks on Cryogenic Transfer Line was started as follows:
E4-F1 leak at F1-end bellows was verified and to be fixed once CHL NRL plant is off.
D1-D2 leak checking is in progress.
D3-D4 leak checking is in progress.
Five IHI cold compressors were removed for inspection and overhaul.
Various FRIG bldg vacuum maintenance is in progress.
Controls Dept is replacing FRIG Front End UPS units with new ones.

Facility Engineering Services
ICW Flush
Started 2/27/06
From Casey’s pond to Proton Site 38 area
Flushed out ~240 pounds of zebra mussel shells
30% complete

Lighting
Pbar completed
MI-8 started
Linac started

CUB – This weekend, 3/4/06
Feeder 42 work
Water system maintenance on the LCW & Chilled water pumps and main tower and tank equipment
All Cooling systems down at 3/3/06 – 4:30 PM to 3/5/06 – 8 PM
No nitrogen from CUB starting 3/4/06

LCW
TeV LCW off starting on 3/4/06 at 7 AM (A0 will stay up)
Booster LCW off starting on 3/6/06 for 6 weeks

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.


last modified 03/03/2006    email Fermilab